OpenStudy (dimpalk104):

Please help!!! How many grams of each element are present in the 8.7 g sample of iron II sulfide, FeS: Mass of Fe in the sample: ?? grams Mass of S in the sample: ?? grams

OpenStudy (greatlife44):

you could calculate the molar mass of each sample first

OpenStudy (dimpalk104):

Fe- 55.85 S-32 FeS-87.85 @greatlife44

OpenStudy (greatlife44):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Dimpalk104 Fe- 55.85 S-32 FeS-87.85 @greatlife44 \(\color{#0cbb34}{\text{End of Quote}}\) Yeah, I would calculate the % of each first like this: we take the mass of sulfur and put it over the total mass of the compound to get the % FeS IS 36% Sulfur so we subtract that by 100 and get 64% Fe by mass \[%Sulfur \frac{ 32 }{ 87.85 } = 36% sulfur S ; 64% Fe \] then we take the % and multiply it by the total mass of the sample 0.64*8.75g = 5.6g Iron 0.36*8.75g =3.15g Sulfur

OpenStudy (michele_laino):

Hint: the mass of \(\text{FeS}\) is: \(55.85+32=87.85\) so we have: \(8.7/87.85=0.099\) mol of \(\text{FeS}\), and \(0.099\) mol of \(\text{FeS}\) contains 0.099 mol of \(\text{Fe}\), and 0.099 mol of \(\text{S}\) so we have: \(0.099⋅55.85=...?\) grams of \(\text{Fe}\), and \(0.099⋅32=...?\) grams of \(\text{S}\)